OpenStudy (anonymous):

A square of side x is inscribed in a circle. Express the area of the circle as a function of x.

OpenStudy (anonymous):

If a square is inscribed in a circle then its diagonal will be the diameter of the circle here, side of square = x therefore diagonal = \[\sqrt{2}x\] = diameter of circle area of circle =[ \[\pi\]*(diameter)^2]/4 = pi*x^2/2

OpenStudy (anonymous):

Thank you!